A rectangle with height and width equal to 3 and 25 respectively, is drawn on a checkered paper. Bazil paints a random horizontal 1×2 rectangle, and Peter paints a random vertical 2×1rectangle (each rectangle consists of 2 sells). Find the probability that at least one of the cells is painted twice. Express the answer in percent, and round to the nearest integer.
1
Expert's answer
2020-12-01T03:28:22-0500

As we can see, it is possible to paint 25 different horizontal "1\times2" rectangles. In a similar way we conclude that on rectangle "3\times 19" it is possible to paint "3 \cdot 24 =72" horizontal "1\times2" rectangles and "2 \cdot 25=50" different vertical "2\times1" rectangles. We have "3600" possible combinations in total.

On "2\times2" square it is possible to paint 4 different intersections of vertical and horizontal rectangles. We can consider "2 \cdot 24=48" different "2\times2" squares on the rectangle. Thus, we obtain "48 \cdot 4=192" different combination, when two rectangles intersect each other. I.e., there will be cells painted twice.

Thus, the probability that one cell is painted twice is: "p=\frac{192}{3600}=0.05 = 5\%"